13) An airline estimates that 95% of people booked on their flights actually show up. If the airline books 77 people on a flight for which the maximum number is 75, what is the probability that the number of people who show up will exceed the capacity of the plane?

A) 0.078

B) 0.253

C) 0.019

D) 0.097
Answer #1

it is binomial distribution :

x = no. of people who show up

n=77

p=0.95

1-p=0.05

probability that the number of people who show up will exceed the capacity of the plane = P(x>75)

P(x>75) = P(76) + P(77)

= 77C76 * 0.95^76 * 0.05^1 + 77C77 * 0.95^77 * 0.05^0

P(x>75) = 0.097

ANSWER : (D) 0.097

(please UPVOTE)